Using the formula for computing future value using compound interest determine the value of an account at the end of 6 years if a principal amount of $12,000 is deposited to an account at an annual interest rate of 6% and the interest is compounded quarterly.


1
Expert's answer
2021-04-27T00:44:03-0400

"FV=PV(1+i)^n"

"i=6\%/4=1.5\%"

"n=4\times6=24"

"FV=12000(1+1.5/100)^{24}=12000\times1.015^{24}=\$17,154.03"